### v1

# CHOICE: Coordinated Human-Object Interaction in Cluttered Environments for Pick-and-Place Actions

Abstract.
Animating human-scene interactions such as pick-and-place tasks in cluttered, complex layouts is a challenging task, with objects of a wide variation of geometries and articulation under scenarios with various obstacles. The main difficulty lies in the sparsity of the motion data compared to the wide variation of the objects and environments as well as the poor availability of transition motions between different tasks, increasing the complexity of the generalization to arbitrary conditions. To cope with this issue, we develop a system that tackles the interaction synthesis problem as a hierarchical goal-driven task. Firstly, we develop a bimanual scheduler that plans a set of keyframes for simultaneously controlling the two hands to efficiently achieve the pick-and-place task from an abstract goal signal such as the target object selected by the user. Next, we develop a neural implicit planner that generates guidance hand trajectories under diverse object shape/types and obstacle layouts. Finally, we propose a linear dynamic model for our DeepPhase controller that incorporates a Kalman filter to enable smooth transitions in the frequency domain, resulting in a more realistic and effective multi-objective control of the character. Our system can produce a wide range of natural pick-and-place movements with respect to the geometry of objects, the articulation of containers and the layout of the objects in the scene.
character animation, scene interaction, motion capture, motion control, implicit representation, deep learning
journal: TOGccs: Computing methodologies Procedural animationccs: Computing methodologies Motion captureccs: Computing methodologies Motion path planning
Figure 1. A selection of multiple interaction tasks in diverse cluttered scenes performed with necessary hand cooperations.
1. Introduction
Controlling characters to perform pick-and-place motions in cluttered environments with obstacles, such as furniture, has significant applications in video games, virtual reality (VR), and robotics. These environments, often found in household chores, can be surprisingly complex and challenging to automate compared to most paid day jobs. The intricacy of the environment escalates due to diverse object shapes, functionalities, and scene layouts with hierarchical levels of objects and varied spatial relations. Planning viable collision-free trajectories for interacting with household objects, like those found in a real-world kitchen with various surface types and tightly packed items, presents an immense challenge. Synthesizing motions for a full-body character that can effectively navigate and interact in such environments is a fundamental yet crucial issue that requires a well-structured approach.
Existing methods that generate human-object interactions in an indoor scene operate in simplified environments. For example, when grasping and manipulating light weight objects, the target object is often located in an open space with no obstacles, where the character simply needs to step forward and grasp it  (Wu et al., 2022; Ghosh et al., 2023). In reality, however, a milk bottle could be contained in a closed refrigerator that is full of items, or a bowl may be contained in a drawer together with other dishes. A successful interaction with the target object become a multi-step process. First, the character needs to prepare an open environment, by opening the refrigerator or pulling out the drawer. Next, they need to clear a collision-free path for the hand to reach the target object, by moving or even taking out nearby obstacles. Finally, the target object may need to be taken out through a narrow opening without colliding with other objects or the environment. Throughout the entire process, the person will coordinate all parts of the body for stepping, opening and grasping tasks according to the layout of the obstacles: for example, a person might side-step and open the cabinet with one hand while reaching out for the target object with the other hand. Although the whole process involves a series of discrete subtasks, the motion is planned such that the entire process is efficient and transitions are smooth.
To tackle these challenges, we propose an integrated system composed of three key modules: a neural implicit trajectory planner that plans collision-free wrist motions, a bimanual scheduler that provides a set of subgoals in terms of keyframes for achieving the goal, and a real-time character controller that synthesizes the motion to reach the planned trajectories and subgoals with smoothness in the DeepPhase (Starke et al., 2022) latent space. Our interaction system approaches the interaction synthesis problem as a hierarchical goal-driven task. Given the complexity of directly dealing with the abstract goal control signal, which only specifies the target object and action, we break down these abstract signals into three-key-joint goal sequences using the Bimanual Scheduler and Implicit Neural Trajectory Planner. We carefully design the inputs and outputs of each module to progressively transit the abstract and sparse goal towards temporally dense and full-body level motion.
To construct a valid 6D trajectory of an object to be grasped and picked out/put in a cluttered environment with multiple obstacles, we develop a novel data-driven implicit neural representation from sparse motion capture dataset. Our approach is inspired by the recent implicit neural representations of 3D shapes (Park et al., 2019; Marschner et al., 2023), scenes (Mildenhall et al., 2020), and navigation (Li et al., 2021): Their success inspires us to develop a unified representation that describes the spatial-temporal relationship between objects in the environment and the grasping hand. The grasping and placing trajectories from our extensively captured motion dataset could be viewed as an an-isotropic level set propagation process. Using these data as training samples, and SDF of the objects and obstacles as the additional inputs, we construct an implicit time-of-arrival field whose gradient represents the trajectory of the hand to reach the target object at any point in the space. During inference, our implicit field robustly plans the trajectory functioned on unseen objects and furniture placements.
As our motions involve a series of discrete subtasks by different parts of the body, a naive controller can result in unsmooth and unrealistic coordination and transitions when the interaction goal switches. Moreover, subtle and complex locomotion, such as side-stepping and body twisting, are coupled simultaneously with upper-body interactions. Reproducing such motion realism from demonstration is challenging. To cope with this issue, we build our interaction controller upon the DeepPhase model (Starke et al., 2022), which has demonstrated high motion quality for acyclic motions like dancing. It is well-known that the frequency domain signals can retain the motion details while simplifying the motion representation. We found that the phase provides a compact and efficient low-dimensional feature space to help formulate the full-body motion, serving as a suitable goal feature for control. We then simplify the control law, making it easier to learn after abstracting the motion into our designed DeepPhase state space. Furthermore, we propose a linear dynamic model for our DeepPhase  (Starke et al., 2022) controller that incorporates a Kalman filter to enable smooth state transitions in the frequency domain. This approach further bridges the sparse interaction instructions to continuous-time goals ahead, resulting in a more realistic and effective control of the character while always following the interaction constraints.
To also plan the sequence of motions that are required for reaching the target object, we propose a scheduler for bi-manual coordination. By integrating all our proposed methods, we create a comprehensive interaction system.
We demonstrate our system’s ability to synthesize realistic pick-and-place interactions in highly complex kitchen environments (see Fig. 1). The system can produce subtle movements such as stepping, turning, and manipulation co-temporarily done by two hands in quick response to the user instructions; the system can generalize to diverse scene layouts and shapes.
The key contributions of this work can be summarized as follows:
  • A systematic method for synthesizing long-term interaction in complex environments, hierarchically integrating three innovative sub-modules with significant enhancements individually, verified through quantitative and qualitative experiments,
  • an extensive motion dataset that includes precise fully-body manipulation including the motions of fingers and multiple types of objects, in contextual lifelike indoor scenes, which consists of motions such as taking out/putting in objects from/into containers, manipulation such as pouring, taking off the cap, etc.,
  • a neural implicit trajectory planner that generates realistic and plausible grasping/placing hand trajectories according to the object type, location and layout of the environment,
  • a DeepPhase-based interaction controller that robustly tracks goals of two-hands and the hip and
  • a bimanual scheduler that produces a series of keyframes for the hip and the wrist for pick-and-place tasks of objects within/to be put in a various containers, including shelves, drawers and cabinets.
2. Related Works
In this section, we first review learning-based motion generation works, especially the RL-based and auto-regressive methods that have the ability for online character-scene interaction. We then review research that develops implicit neural representations for the purpose of environmental perceptions and motion planning.
Motion Synthesis of Character-scene Interaction
The field of motion synthesis has recently seen rapid advancements, utilizing both data-driven and physics-based methodologies (Peng et al., 2021, 2022; Tessler et al., 2024). In the realm of human-scene interaction, kinematics-based humanoid controllers have proven effective in executing interactive motions such as sitting, opening doors, and carrying objects (Starke et al., 2019; Zhang et al., 2022). By incorporating generative frameworks such as conditional Variational Autoencoders (cVAE) (Hassan et al., 2021), Diffusion models (Pi et al., 2023; Cen et al., 2024), or by further learning a Reinforcement Learning (RL) policy to control the VAE’s learned latent space (Zhao et al., 2023a; Luo et al., 2023), a more diverse range of interaction behaviors can be synthesized. In the application of physics simulation, diverse interaction motions such as catching and carrying, coupled with locomotion skills, have been realized by distilling a variety of expert demonstrations and RL-based control with visual input (Merel et al., 2020). More recent progress involves incorporating adversarial training in imitation learning to synthesize realistic interactions (Hassan et al., 2023). While these approaches yield impressively realistic animations, they are often constrained by the type of objects involved or the complexity of the scene.
Enhanced by the construction (Hassan et al., 2019) or capture (Araujo et al., 2023) of real-life scene and motion pairs, human interactions can be synthesized in more clustered and complex scenarios(Li et al., 2023a). By extending the PROX dataset (Hassan et al., 2019) with semantic labels and scene segmentation, interaction poses can be conveniently generated through semantic control (Zhao et al., 2022). To compensate for datasets that lack motion-scene pairs, LAMA (Lee and Joo, 2023) and DIMOS (Zhao et al., 2023b) are introduced to fulfill appropriate motions in the desired scene interaction by incorporating RL. Although these methods can adapt to clustered motions, most interactions are limited to static scenes, lacking details of dynamic objects and finger motions.
To produce fine-grained finger motions for object interaction, a wide range of methods have been explored, including the introduction of canonicalized contact representations (Zheng et al., 2023), grasp fields (Karunratanakul et al., 2020), and hand-object spatial representations (Zhang et al., 2021). Leveraging a physics engine, contact sampling (Ye and Liu, 2012), and DRL-based control (Christen et al., 2022; Zhang et al., 2023; Hu et al., 2024) are utilized to synthesize physically plausible manipulations. While these methods focus on isolated hand-object interactions, recent approaches have learned to synthesize full-body motions driven by dynamic object movements (Li et al., 2023b, a). Incorporating part-wise motion priors (PMP) (Bae et al., 2023) also allows for the physical assembly of hand and body motions that adapt to complex object movements. More recently, high-quality full-body object manipulation motion datasets have been acquired (Taheri et al., 2020; Bhatnagar et al., 2022; Fan et al., 2023), providing opportunities to explore full-body grasping pose synthesis (Xu et al., 2023) and motion control (Taheri et al., 2022; Wu et al., 2022; Ghosh et al., 2023; Braun et al., 2024). While GOAL (Taheri et al., 2022), GRIP (Taheri et al., 2024), and SAGA (Wu et al., 2022) concentrate on synthesizing full-body approaching and grasping motions, human-object interactions are further enhanced by conditioning on human intention (Ghosh et al., 2023) or physics-based hierarchical control (Braun et al., 2024; Luo et al., 2024). However, existing full-body object interaction works are still limited to relatively short motion sequences and monotonous scenes. Drawing from our interaction dataset, we propose a method for synthesizing full-body motion that engages with a variety of scenarios, featuring an array of furniture and everyday objects. Furthermore, we demonstrate that our framework is capable of executing a series of complex tasks involving object and scene interactions, all under interactive user control.
Our control framework design is inspired by a stream of research that has been proposed to discover the periodic motion feature. This includes the periodical feature describing the feet-ground contacting and controlling the character using the phase in the frequency domain (Holden et al., 2017; Zhang et al., 2018; Starke et al., 2021). DeepPhase (Starke et al., 2022) has further proposed the Periodical Auto-Encoder (PAE) that computes a low-dimensional representation of the spatial-temporal body motion. By running a non-linear transformation and FFT, PAE can encode the full-body joint-space motion into several channels of periodic latent vectors, forming a phase manifold in the frequency domain.
The advantage of the approach is that it can spatio-temporarily align a wide range of cyclic and acyclic movements with different hand and leg movements to form a motion manifold in the frequency domain, and then allowing detailed auto-regressive control of the body through trajectories projected on the ground. Such a characteristic is attractive for our pick-and-place task that requires the combination of complex leg movements, e.g., side-stepping and turning, and co-temporal hand movements, e.g., opening the cupboard on one hand while grasping an object with the other hand, that also requires position-based control for guiding the end effectors to the reach the target while avoiding obstacles in the environment.
Implicit Neural Representation for Motion Planning
The primary concept of using implicit neural representation (INR) methods for 3D geometry involves learning a parametric function via a neural network to represent an objective, such as the signed distance for the query coordinates (Park et al., 2019; Marschner et al., 2023) or the radiance field for the query view (Mildenhall et al., 2020). Following its success in shape reconstruction tasks, recent studies have also explored the use of INR-based functions for motion-related activities. By constructing the affordance function to determine a potential grasping pose, implicit neural methods have been proposed to learn a field in SE(3) space for the gripper (Simeonov et al., 2022; Manuelli et al., 2019; Chen et al., 2022) or for the finger-hand mesh (Karunratanakul et al., 2020; Yang et al., 2022). In the realm of navigation, Li et al. (2021) have utilized an implicit environment field that encodes the maze map and predicts the reaching distance to the target position. Additionally, Camps et al. (2022) have developed an online deep SDF of the surrounding environment, and NTFields (Ni and Qureshi, 2022) navigates the robot through a time field. Because of its continuity in the latent space, INR naturally extends to feasible grasping for deformed object shapes and navigates the agent using a distance field that fits a different scene.
Most of the current trajectory planning methods are designed for robotic applications, and the trajectory optimization in the canonical space is usually conducted after configuring their potential field (Simeonov et al., 2022; Wang et al., 2024). Similarly, Chen et al. (2022) apply model predictive control to optimize the objective function that incorporates the evaluation from the grasping field or the distance field, respectively. However, kinematic-based motion generation for virtual characters can often yield more human-like motion patterns. These could include subtle nuances like slightly lifting an object when carrying it out from an upper shelf, or keeping the in-hand object moving with redundant space to potential obstacles while approaching more straightly if the hand is empty. Performing such motion patterns is opposed to merely achieving the objective with the least control cost of the robot arm. Therefore, the direct application of trajectory planning will fail to execute some realistic motion patterns that require more control energy. To address this, our proposed implicit neural trajectory planner eliminates the need for an extra stage of adapting trajectory planning algorithms. Instead, it directly learns the realistic wrist trajectory conditioned on the state and environment from the MoCap dataset.
Figure 2. Overview of our CHOICE System: Perceiving the involved objects around the clicking goal object with the action instruction from the keyboard, our system arrange tasks and motions to the empty hands, and match bi-manual goals based on the character state and the goal tasks. The matched hand goal priors are then re-planned by our trajectory planning sub-system to generate a trajectory of manipulation goals that fit the runtime environment. After planning a set of navigation goals together with the manipulation trajectory, our goal coordination arranges the goal of the key-joints and a goal phase prior to the current character. The second stage of our system is the goal-driven motion controller, it auto-regressively updates the character motion towards the goal of keyjoints, together with a loop of the Kalman filter which corrects and sets the goal phase features.
3. System Overview
The structure of our real-time character control system is shown in Fig. 2. We split the interaction synthesis task into three major sub-tasks:
  • hand-object trajectory planning, managed by the trajectory planning sub-system (brown) shown in  Fig. 2 and § 4.2;
  • bimanual interaction scheduling, managed by the goal matching sub-system (grey) and the goal coordination (cyan) shown in Fig. 2 and § 6.3; and
  • fully-body motion control, managed by the goal-driven control sub-system (blue) shown in Fig. 2 and § 5.
When the user clicks the object to pick or place, the path to walk towards the object is first planned, then the series of keyframes for the key joints (left/right wrist and hip) based on contacts are scheduled and finally, collision-free trajectories of the arms to conduct the motion is computed. The DeepPhase-based controller then synthesizes the motion to follow the scheduled sub-goals and planned trajectories in real-time in the phase manifold. The major mathematical variables involved in these three sub-systems are summarized in Table. 1
Table 1. Overview of Major Mathematical Variables
Sec. Var. Description
4. Traj. Plan $\mathbf{x}$ Query position
$\mathbf{c}$ Interaction conditions
$\mathbf{z}$ Latent vector of the auto-decoder
$D_{t}$ Distance field of the target
$D_{o}$ Distance field of obstacles
$D_{toa}$ Time-of-arrival field of a hand trajectory
$\mathbf{x}_{prior}$ Accessiable part of $D_{toa}$ in testing
$\tau_{prior}$ A matched trajectory prior as in § 6.1
$\mathbf{x}_{s}$ The predicted trajectory starting position
5.1 Features $\mathbf{T}_{t}^{key}$ 2s trajectory of a key joint pivot on time $t$
$\mathbf{P}_{t}$ Phase feature vector pivoted on time $t$
$g$ Timestamp of the next goal ahead
$\mathbf{G}^{keyJ}$ Goal transformation of all the keyjoints
$\mathbf{T}_{t}^{ego}$ Ego-centric keyjoint trajectory prediction
$\mathbf{T}_{t}^{inv}$ Goal-centric keyjoint trajectory prediction
5.2 Controller $\mathbf{X}_{t}^{p}$ Phase feature pivoted at $t$ as state
$\mathbf{Y}_{t}$ Body pose prediction at $t$
$\mathbf{U}$ Network’s control law of phase feature
$\mathbf{G}^{p}$ Goal state of the phase feature
$\mathbf{P}$ Covariance matrix of KF estimation
$\mathbf{R}_{m}$ Covariance matrix of a measurement
6. Scheduler $\bm{p}$ Character’s current pose when matching
$\bm{N}$ Navigation matching feature vector
$M$ Hand-object interaction matching function
4. Neural Implicit Trajectory Planner
In this section, we present the neural implicit trajectory planner designed for the sub-task of generating precise hand grasping or placement of lightweight objects in cluttered scenes with functional furniture (see Fig. 2, trajectory planning sub-system in brown). Implicit representation has exhibited remarkable success in learning human face and body shapes, as well as 3D objects. The motivation behind employing the implicit representation for trajectory planning stems from its exceptional ability to generalize from limited training data. The pick-and-place motion involves intricate object handling, which encompasses various conditions such as diverse geometries, shapes, and articulation of the container, as well as object locations within the container. Additionally, the trajectory planning is influenced by the presence of obstacles of different shapes at various positions along the path, along with the position and orientation of the character executing the task. By leveraging the implicit representations, we can harness its capacity to handle the complexity and variability of these conditions, empowering the controller to devise trajectories that seamlessly adapt to a wide range of scenarios encountered during the pick-and-place task.
We introduce an implicit trajectory planner based on an auto-decoder structure, where the trajectory is inferred by optimizing a time-of-arrival field that represents the trajectory to reach or place the object from all directions. This approach draws inspiration from DeepSDF (Park et al., 2019), where an auto-decoder structure is applied to learn spatial implicit fields, providing a compact and continuous latent space that bridges the motion and the scene.
Although DeepSDF can result in over-smoothed surface reconstructions, losing some high-frequency details, this drawback is not critical for our trajectory planning. To generate smooth and collision-free trajectories, capturing the general shape and position of obstacles is more important than preserving fine surface details, and the region-of-interest for obstacle avoidance is distributed away from the obstacle surfaces. Indeed, current state-of-the-art trajectory planners (Wang et al., 2022; Zhou et al., 2019; Sundaralingam et al., 2023) mostly simplify the detailed surface of obstacles with convex primitives.
We first introduce a new field representation that models the hand movements jointly with the geometric information from the environment in § 4.1. Then we describe our auto-decoder architecture (Fig. 3), and how it is trained to synthesize an environment-aware hand trajectory in § 4.2.
4.1. Joint Neural Representation of the Hand Trajectory and the Environment
We describe the 3D space $E$ using features that contribute to the trajectory planning of the hand approaching and leaving. The backbone of our system is a function $\mathbf{f}$ that maps each local coordinate $\mathbf{x}\in R^{3}$ to a feature vector $\mathbf{D}$ as follows:
(1) $\mathbf{f}:\mathbb{R}^{3}\rightarrow\mathbb{R}^{3},\mathbf{f}(\mathbf{x})=\mathbf{D}(\mathbf{x})=(D_{t}(\mathbf{x}),D_{o}(\mathbf{x}),D_{toa}(\mathbf{x}))$
where $D_{t}(\mathbf{x})$ is the inverse unsigned distance of each position $\mathbf{x}$ to the target object surface $M_{t}$, $D_{o}(\mathbf{x})$ is similarly the inverse unsigned distance to the set of obstacles $M_{O}:=\{M_{O_{1}},M_{O_{2}},...,M_{O_{n}}\}$, and $D_{toa}(\mathbf{x})$ represents the time to be travelled by the hand between $\mathbf{x}$ for reaching the object: we now describe the details of each term below.
$D_{t}(\mathbf{x})$ is computed using the fast marching method (FMM) (Tsitsiklis, 1995; Sethian, 1996) following the Eikonal equation
(2) $\|\nabla d(\mathbf{x})\|=\begin{cases}1&\text{if }\mathbf{x}\ outside\ M_{t}\cup M_{o}\\ \infty&\text{if }\mathbf{x}\ inside\ M_{t}\cup M_{o}\end{cases}$
with the boundary condition $d(\mathbf{v}_{i})=0$ for all vertices $\mathbf{v}_{i}$ on $M_{t}$. $D_{t}(\mathbf{x})$ then forms a spatially continuous function as follows:
(3) $D_{t}(\mathbf{x})=1/\max(d(\mathbf{x},M_{t}),\epsilon)$
where the reciprocal form makes the field continuous on the boundary near the obstacles, and clipping the unsigned distance makes the field continuous on the boundary near the target object’s surface.
$D_{o}(\mathbf{x})$, the inverse distance field of the obstacle, is similarly formed as follows:
(4) $D_{o}(\mathbf{x})=1/\min\{\,\max(d(x,M_{O_{i}}),\epsilon)),i=1,2,...,n\}.$
$D_{toa}(\mathbf{x})$, the time of arrival field, represents the 6D trajectory of the hand to reach/leave the mesh surface; the field is learned from each motion capture data and is extrapolated to the volume outside the trajectory and can further generalize to unseen scenes. We first describe how we compute $D_{toa}$ for each training data for reaching out for a target object. We generate a potential field $\phi(\mathbf{x})$ such that its value on every point $\mathbf{x}$ on the hand trajectory $\tau$ represents the remaining time for the wrist to move from the current point to $\mathbf{x}_{goal}$ where the hand first contacts the object.
We evolve the time-of-arrival of the demonstration trajectory $\tau$ to the 3D space using the Gaussian probability density function (PDF), which decreases the speed value of each grid according to its distance to the trajectory $\tau$. Setting the initial value as $\phi(\mathbf{x}_{goal})=0$, $\phi$ is computed by solving the following Eikonal equation:
(5) $\|\nabla\phi(\mathbf{x})\|=\begin{cases}\frac{1}{||\mathbf{v}(\mathbf{x}_{n})||\cdot PDF(d(\mathbf{x},\tau),\sigma)}&\text{if }\mathbf{x}\ outside\ M_{t}\cup M_{O}\\ \infty&\text{if }\mathbf{x}\ inside\ M_{t}\cup M_{O}\end{cases}$
where $\mathbf{x}_{n}$ is the nearest position on the trajectory $\tau$ from $\mathbf{x}$ and the hand speed at $\mathbf{x}_{n}$ $||\mathbf{v}(x_{n})||$ is set as the peak value of the Gaussian PDF. Finally, $D_{toa}$ is computed as an inverse of the time-of-arrival as below:
(6) $D_{toa}(\mathbf{x})=1/\max(\phi(\mathbf{x}),\epsilon))$
which ensures the continuity of the field.
Consequently, the three fields formed by $\mathbf{f}$ describe three different but highly related types of representation under the same environment: Querying each position $\mathbf{x}$, we can acquire its reaching distance to the goal object, the reaching distance to the nearest obstacle, and an arrival time to the contact frame. The three fields are set as the outputs of an implicit neural function $\mathbf{f}_{\theta}$ which is applied to address the scene understanding and trajectory planning tasks in a combined fashion as described next in § 4.2.
Figure 3. Implicit neural trajectory planner: The scene visualisation of the three fields uses the green color to show its value, deeper green represents a lower distance, and the blue color highlights the region of infinity distance. The right-side images give the 2D slices at the teapot height, where the zero-level set was shown in orange curves. Under a test scene, $\mathbf{z}$ was optimized to reconstruct the known part of the output, which was encircled by the blue rectangles. The dashed blue rectangular illustrates the pre-known part of the time-of-arrival field.
4.2. Implicit Neural Trajectory Planning Architecture
In this section, we form an autodecoder architecture as shown in Fig. 3, which we apply to compute the trajectory to reach or release an object in the environment. This is done by optimizing the latent code $\mathbf{z}$ and computing the time-of-arrival field $D_{toa}$. To achieve this task, we first train the autodecoder using sample fields computed from the motion capture data described in § 7. During runtime, using the trained autodecoder, we optimize the latent code $\mathbf{z}$ such that the loss function that evaluates the L1 loss of the fields and other regularization losses is minimized.
Architecture:
As shown in Fig. 3, the autodecoder receives the latent code $\mathbf{z}$, the condition vector $\mathbf{c}$ and the current position $\mathbf{x}$ as the inputs and outputs the three field values $D_{t}(\mathbf{x})$, $D_{o}(\mathbf{x})$ and $D_{toa}(\mathbf{x})$. The latent code $\mathbf{z}\in\mathbb{R}^{128}$ encodes the environment, the object geometry as well as the hand trajectory, and is computed by optimization both during training and testing. The condition vector $\mathbf{c}\in\mathbb{R}^{4}$ describes the hand state tokens from the real-time controller composed of the height of the goal object, the label of the currently controlled hand (left or right), the goal action label (grasping or placing), and the label of the current trajectory segment (approaching or leaving). The composition of this information provides the semantic and spatial information that influences the trajectory. Together with the environment geometry, most indoor pick-and-place motion conditions can be described.
Training:
We concatenate the latent code $\mathbf{z}$ with the condition $\mathbf{c}$ and the querying position $\mathbf{x}$, and optimize the network parameter $\theta$ and $\mathbf{z}_{i}$, which represents the $i$-th captured scene. At the training stage, the system learns the mapping from the condition $\mathbf{c}$ and query position $\mathbf{x}$ to the $\{D_{t},D_{o},D_{toa}\}$ features:
(7) $\mathcal{L}_{\text{Rec}}=\sum_{\mathbf{x}\in E}\bigg{(}$ $\left\|f_{\theta}(\mathbf{x},\mathbf{z};\mathbf{c})_{0}-D_{t}(\mathbf{x})\right\|_{1}+\left\|f_{\theta}(\mathbf{x},\mathbf{z};\mathbf{c})_{1}-D_{o}(\mathbf{x})\right\|_{1}+$
$\left\|f_{\theta}(\mathbf{x},\mathbf{z};\mathbf{c})_{2}-D_{toa}(\mathbf{x})\right\|^{2}_{2}\bigg{)}.$
Note $L_{2}$ loss is used for $D_{toa}$ while $L_{1}$ loss is used for $D_{t}$ and $D_{o}$: This is because $L_{2}$ loss is more sensitive to the nonuniform distribution caused by the anisotropic propagation of $D_{toa}$ and thus suitable.
We then train the network with all $N$ approaching/leaving scenes by following the maximum log posterior with respect to the latent code $\mathbf{z}$ and the network parameters $\theta$ as (Park et al., 2019):
(8) $\underset{\theta,\left\{\mathbf{z}_{i}\right\}_{i=1}^{N}}{\arg\min}\sum_{i=1}^{N}\left(\mathcal{L}_{\text{Rec}}+\frac{1}{\sigma^{2}}\left\|\mathbf{z}_{i}\right\|_{2}^{2}\right)$
where $\sigma=0.01$ as in DeepSDF (Park et al., 2019).
Runtime:
We estimate the hand motion by reconstructing the $D_{toa}$ field whose gradient corresponds to the hand trajectory; this is done by optimizing $\mathbf{z}$ such that the field reconstructed through $f_{\theta}$ matches the partially observable field values:
(9) $\hat{\mathbf{z}}=\underset{\mathbf{z}}{\mathrm{argmin}}\Bigg{(}$ $\sum_{\mathbf{x}\in E_{t}}\left(\|f_{\theta}(\mathbf{x,z};\mathbf{c})_{0}-D_{t}(\mathbf{x})\|_{1}+\|f_{\theta}(\mathbf{x,z};\mathbf{c})_{1}-D_{o}(\mathbf{x})\|_{1}\right)$
$+\gamma\sum_{\mathbf{x}\in\mathbf{x}_{prior}}\|f_{\theta}(\mathbf{x,z};\mathbf{c})_{2}-D_{toa}(\mathbf{x})\|^{2}_{2}+\frac{1}{\sigma^{2}}\|\mathbf{z}\|^{2}_{2}\Bigg{)}$
where $\mathbf{x}_{prior}=\{\mathbf{x}_{O},\mathbf{x}_{T}\}$ represents the set of locations that $D_{toa}$ is known, including
  • $\mathbf{x}_{O}$: $\mathbf{x}\ inside\ M_{O}$, such positions shall never be approached by the hand, thereby $D_{toa}=0$;
  • $\mathbf{x}_{T}$: The set of $\mathbf{x}$ when the grasping/placing is in action. Given $\mathbf{c}$ and $D_{t}$, a grasping/placing motion clip is matched in the database, whose wrist trajectory is defined by $\tau_{prior}$. We include $\mathbf{x}\ \in\ \tau_{prior}$ where $d(\mathbf{x},M_{t})<5cm$ in $\mathbf{x}_{T}$,
$E_{t}$ is an object-centric grid aligned with surrounding furniture, and $\gamma$ re-weights the loss of $D_{toa}(\mathbf{x})$ to balance it with the other two losses according to the number of samples in the grid: $\gamma=N(E_{t})/N(\mathbf{x}_{prior})$. The optimized latent variable $\hat{\mathbf{z}}$ and the decoder $f_{\theta}$ together form an implicit neural function that generates a wrist trajectory, taking into account its spatial relationship with the target object and surrounding obstacles.
During the goal approaching, the wrist moves along the descending gradient of the estimated time-of-arrival field $\hat{\phi}=1/\hat{D}_{toa}$, and we can acquire the wrist velocity $\mathbf{v}$ at each $\mathbf{x}$ by:
(10) $\mathbf{v}(\mathbf{x})=\nabla\hat{\phi}(\mathbf{x})^{\circ-1},$
where ∘-1 is the Hadmard inverse operation. The velocity direction is determined by the inverse Hadamard product with the coordinate vector. Thereby we can update the wrist position with $\mathbf{v}(\mathbf{x})$. Because the goal grasping or placing position corresponds to a sink point of $\hat{\phi}$ where the gradient is zero, an integration starting from this point is unstable. We thus find a position between the wrist position $\mathbf{w}$ at the planning frame and the object to start the trajectory integration:
(11) $\mathbf{x}_{s}=\underset{\mathbf{x}}{\mathrm{argmin}}\{-\hat{\phi}(\mathbf{x})+d(\mathbf{x},\mathbf{w})/\bar{v}\},\ s.t.\ \ \hat{\phi}(\mathbf{x})\neq+\infty$
where $\bar{v}$ is the average wrist speed in a scalar. We then integrate from $\mathbf{x}_{s}$ by gradient descent to compute the positional elements of the newly planned tangent approaching/leaving trajectories of the wrist $\mathbf{t}^{\prime}$, where the order of the leaving points are then reversed.
The 6D trajectory of the wrist $\tau_{plan}$ that follows the planned tangent trajectory $\mathbf{t}^{\prime}$ is computed by utilizing the matched wrist motion $\tau_{prior}$, and then transferring the orientation $\mathbf{R}_{1:m}$ from $\tau_{prior}$ to that of the newly planned trajectory. Specifically, the orientation of the $(m-i)$-th frame from $\tau_{prior}$ is added to the planning frame of $\mathbf{x}_{n-i}$, where $m,n$ are the frames synchronizing $\tau_{prior}$ and $\tau_{plan}$ based on the contact - the first contact frame for picking and release frame for placing. Mathematically, the $3\times 3$ rotation matrix $\mathbf{R}^{\prime}_{1:m}$ of the $i$-th frame of the new trajectory $\tau_{prior}$ is computed by the following process:
(12) $\mathbf{v}$ $=\mathbf{R}_{m-i}(:,j)\times\mathbf{t}_{m-i};\quad\theta_{j}=\arccos(\mathbf{R}_{m-i}(:,j)\cdot\mathbf{t}_{m-i})$
$\mathbf{R}^{j}_{align}=\mathbf{I}+[\mathbf{v}]_{\times}\sin(\theta_{j})+[\mathbf{v}]_{\times}^{2}(1-\cos(\theta_{j}))$
$\mathbf{R}^{\prime}_{n-i}(:,j)=\mathbf{R}^{j}_{align}\mathbf{t^{\prime}}_{n-i}\quad\quad(j=0,1,2)$
where $[]_{\times}$ is the cross product operation matrix, $\mathbf{t}_{m-i},\mathbf{t^{\prime}}_{n-i}$ are the tangent vectors on $(m-i)$/$(n-i)$-th frames in $\tau_{prior},\tau_{plan}$. Furthermore, we linearly blend the final approaching frames on $\tau_{plan}$ towards $\tau_{prior}$ to ensure the object touching/releasing constraint is preserved.
In summary, the matched prior determines the hand-object contact and realistic wrist orientations, and our runtime optimization identifies a latent code optimally tailored to the test environment (blue boxes in Fig. 3), which is mapped to the time-of-arrival field that reliably extracts a collision-free positional path. Using Eq. 12 a plausible 6D manipulation trajectory is produced.
5. Deep Phase Interaction Controller
Figure 4. Framework of Our Deep phase Interaction Controller: The Kalman filter estimates the target phase correlated to the goal key-joint transformations. The Gating Network compares the feature of key-joint transformations and the phase feature from the current and goal, and after each motion prediction for the next frame, our bi-directional control blends the key-joint transformation prediction, also feed-back the covariance to the Kalman filter based on the displacement of the bi-directional prediction.
In this section, we describe our real-time motion generator that robustly tracks the sparse trajectories provided by the hand trajectory planner in § 4.2 and the root motion planned by a 2D planner (see § 6.2), as Fig. 2, goal-driven motion control sub-system (blue). Our framework follows the DeepPhase controller (Starke et al., 2022) but with extensions to control the hands and root of the body to move towards their goals, such as done in the Neural State Machine (NSM) (Starke et al., 2019). The NSM requires all motions to be phase-labelled based on predefined phases; instead we make use of the phase labels computed by the Periodic Autoencoder (Starke et al., 2022) and develop an algorithm to guide the character towards the goal in the phase space (see § 5.1). Because the phase prediction at the goal is highly uncertain, we develop a scheme based on Kalman filter to continuously improve the prediction over time (see § 5.2).
The system is provided with the goal states of the keyjoints, i.e., root, hip and the two hands for guidance: these are planned by a high level controller that is later described in § 6.
5.1. Real-time Goal Directed Controller
Our system has a Mixture-of-Experts (MoE) architecture that autoregressively predicts the state of the character for real-time control (Zhang et al., 2018; Starke et al., 2019, 2022); it receives the current pose of the character and the control signals, and outputs the pose in the next frame as well as the revised control signals (see (Zhang et al., 2018; Starke et al., 2019) for the details). We give the goal state of some key joints to guide the character towards the keyframe; in addition to the 2D root trajectories on the ground (Starke et al., 2019), the 6D transformations of the hip and left/right wrists are also given for precise pick-and-place motion synthesis. Further, the phase features at the goal states are given as additional control signals. The details of the newly introduced input/output goal features are given below.
Input Features:
Goal-focused Features $\mathbf{G}_{t}^{keyJ}=\{\mathbf{T}_{g}^{hip},\mathbf{T}_{g}^{root},\mathbf{T}_{g}^{left},\mathbf{T}_{g}^{right}\}$ set the goal-related transformation signals of the three key-joints together with the body root projected onto the ground, extracted from the goal time window pivoted on frame $g$, which is the next keyframe after current timestamp $t$. The keyframes are inserted when the following two types of events happen: (1) the action label of the body changes, such as when switching to “walking” from “idling” and (2) when the pick-and-place motion starts, ends and the contact state with the object changes in the middle.
[…]
The Noised Phase Features at the Goal $\mathbf{P}_{g}+\bm{\epsilon}$. The phase features at the goal keyframes are also provided as inputs, which help to align the timing of the training data (Starke et al., 2019), though here we use the phase by the DeepPhase (Starke et al., 2022). These are precomputed for the training data by the PAE and given as the input. Gaussian noise $\bm{\epsilon}\in\mathcal{N}(0,0.1)$ is added to the phase features $\tilde{\mathbf{P}}_{g}=\{\tilde{\mathbf{p}}_{g},\tilde{\mathbf{F}}_{g}\}$ to enhance the model’s robustness to input variations as:
(13) $\tilde{\mathbf{p}}_{g}$ $=\left(\bm{A}_{g}\cdot\cos\left(2\pi\cdot\bm{S}_{g}\right)+\bm{\epsilon}_{r},\bm{A}_{g}\cdot\sin\left(2\pi\cdot\bm{S}_{g}\right)+\bm{\epsilon}_{i}\right)$
$\tilde{\mathbf{F}}_{g}$ $=\bm{F}_{g}+\bm{\epsilon}_{f}$
where $(\bm{A}_{g},\bm{S}_{g},\bm{F}_{g})$ are the phase features at the goal composed of amplitude, phase, and frequency, and $\bm{\epsilon}_{r},\bm{\epsilon}_{i},\bm{\epsilon}_{f}$ are three independent Gaussian noise added on the real/imaginary parts of the phase and the frequency, respectively. And the Gaussian noise added here will be equivalent to a Gaussian noise in the temporal domain after iFFT.
The purpose of the noise is to enhance the network robustness for precise goal tracking, and simultaneously train the network to learn to rectify the biased goal phase features to an unbiased prediction of the goal as well as an unbiased motion updating. Actually, during testing, we cannot get prior knowledge of the phase features at the goal; the network prediction learned from the data will help refine the goal phase features needed for tracking the key-joint goals. We further develop an error-removal mechanism at the control scheme introduced in Section 5.2.
Output Features:
[…]
Goal-centric Key Trajectory Prediction $\mathbf{T}_{t+1}^{inv}=\{\mathbf{T}_{t+1}^{root},\\ \mathbf{T}_{t+1}^{hip},\mathbf{T}_{t+1}^{left},\mathbf{T}_{t+1}^{right}\}$ are the future trajectories of the key joints as the egocentric one, but represented relative to the keyjoint’s coordinate system at the goal: $\mathbf{G}_{t+1}^{keyJ}$. Ideally for each joint, the bidirectional prediction in the transformation matrix representation should follow the constraint:
(14) $\mathbf{T}_{t+1}^{ego}=\mathbf{G}_{t+1}^{keyJ}\mathbf{T}_{t+1}^{inv}.$
although usually there is a discrepancy due to the lower precision of the prediction especially when the goal is far away: we thus blend the two trajectories with increasing weight for the goal-driven trajectory as approaching to the goal.
[…]
Figure 5. The distribution of amplitude and frequency control in $\mathbf{U}$, revealing the natural motion transitions, which exhibit consistent acceleration and deceleration patterns and perform the motion diversity during interactions, follows the zero-mean Gaussian distribution in the frequency-domain latent.
5.2. Goal-driven Control Scheme in Deep Phase Manifold
In this section, we demonstrate that the DeepPhase control can be effectively represented as a linear dynamical system within the DeepPhase latent space. Leveraging this characteristic, we propose a goal-driven controller that enables smooth tracking of goal phase feature vectors in a latent space. The Periodic Auto-encoder of the DeepPhase framework converts high-dimensional nonlinear motion in the canonical space to a low-dimensional representation where linear dynamics can be applied, hence suiting learning long-term interactions. Additionally, we introduce a novel feedback loop by utilizing the network’s predictive capability to estimate the goal phase, and establish a relationship between the measurement covariance in the Kalman filter and the bi-directional prediction bias in the spatial domain. This innovative approach significantly enhances guidance in the frequency domain latent, hereby improve the full-body motion coordinate.
DeepPhase as a Linear Dynamical System
Denoting the current phase states as $\mathbf{X}_{t}^{p}=[\mathbf{S_{t}},\mathbf{A_{t}},\mathbf{F_{t}}]^{T}$, the goal state as $\mathbf{X}_{g}^{p}$ and the body pose as $\mathbf{Y}_{t}$, the DeepPhase (Starke et al., 2022) system dynamics can be written as:
(15) $\mathbf{X}_{t+\Delta t}^{p}=\mathbf{A}\mathbf{X}_{t}^{p}+\mathbf{B}\mathbf{U}$
$\mathbf{Y}_{t+\Delta t}=f_{MoE}(\mathbf{X}_{t+\Delta t}^{p},\mathbf{Y}_{t})$
where $\mathbf{A},\mathbf{B}$ are constant square matrices
(16) $\mathbf{A}$ $=\left[\begin{array}[]{ccc}1&0&\Delta t\\ 0&1&0\\ 0&0&1\end{array}\right],\mathbf{B}=\left[\begin{array}[]{ccc}\beta&0&0\\ 0&1&0\\ 0&0&1\end{array}\right]$
and the network control $\mathbf{U}$ updates the phase variables as:
(17) $\bm{U}$ $=\left[\begin{array}[]{l}\Delta\mathbf{S}\\ \Delta\mathbf{A}\\ \Delta\mathbf{F}\end{array}\right]=f_{MoE}\left(\mathbf{X}_{t}^{p},\mathbf{X}_{g}^{p},\mathbf{Y}_{t}\right)\sim\bm{C}[\mathbf{X}_{g}^{p},\mathbf{X}_{t}^{p}]^{T}$
where $\beta=0.5$ is a blending weight that enforces the phase update to follow the frequency. Among the state variables, the phase is circularly updated and constrained within the range: $\mathbf{S}_{t}\in[0,1]$, the amplitude $\mathbf{A}_{t}\in[0,1]$, and the frequency $\mathbf{F}_{t}\in[0,\mathbf{F}_{m}]$ where $\mathbf{F}_{m}$ is the Nyquist frequency. Since $|\det(A)|=1$ and the system is fully controllable via $\mathbf{U}$, this system exhibits critical stability and with appropriate control inputs, the system state $\mathbf{X}_{t}^{p}$ can be guided to converge toward the desired goal state $\mathbf{X}_{g}^{p}$ in the latent space. Because the controller network is trained to reduce the difference between the goal phase variables $\mathbf{X}_{g}^{p}$ and the current phase variables $\mathbf{X}_{t}^{p}$, we can approximate its output $\bm{U}$ by a linear transformation of the state error, i.e., $C(\mathbf{X}_{g}^{p}-\mathbf{X}_{t}^{p})$, with $\bm{C}$ as a coefficient matrix. Due to the high-order smoothness of natural movements, the updates $\bm{U}$ are small and bounded.
The control signal $\mathbf{U}$ can be statistically modeled as a zero-mean Gaussian distribution (see Fig. 5). This assumption aligns with the characteristics of human motion sequences, which are often represented by Gaussian distributions (Min and Chai, 2012; Mukai and Kuriyama, 2005). As $U$ follows the Gaussian distribution, the process noise (error from directly inferring the next state by $\mathbf{X}_{t+\Delta t}=\mathbf{A}\mathbf{X}_{t}$) and the measurement noise from the controller network also follows the Gaussian distribution. The approximated linear dynamics of the phase state variables, coupled with the Gaussian noise assumption, allow us to use the Kalman filter to achieve stable goal tracking, as described next.
Recursive Goal State Estimation by Kalman Filter
Ensuring precise and continuous target inputs $X_{g}^{p}$ online is crucial for achieving robust goal tracking and precise convergence to the user’s intended interactions. In our system, as presented in § 5.1, we assume that the goal phase is given. However, using the matched goal provides a biased phase prior due to differences in matching, and can only be provided for a short horizon around the action key-frame. To address this challenge and appropriately lead the character, we develop a framework for accurately estimating goal phases aligned with our planned hand trajectory using the Kalman filter. This filter is optimal for estimating the time-varying goal state under linear system dynamics with Gaussian noise. Our Kalman filter incorporates two measurements, namely predictions from the network ($m=pred$) which is unbiased and well suits time-varying goals, and the matched motion data prior ($m=match$) aligned with the same target action. Algorithm 1 outlines our pipeline.
Input: Motion frame $\bm{X}_{t}$, statistical variance $\bar{\bm{\sigma}}$
Output: Goal phase states $\mathbf{G}^{p}$ and Covariance $\mathbf{P}$
Initialize $\mathbf{G}^{p}=0$ and $\mathbf{P}=0$
while  $\mathbf{G}_{t}\neq$ null do
       if MouseClicking then
             Intialize goal $\mathbf{G}_{t}=\{\mathbf{G}_{t}^{p},\mathbf{G}_{t}^{keyJ}\}$
       end if
      Network Inference $\bm{X}_{t+1},\mathbf{T}_{t+1}^{ego^{\prime}},\mathbf{T}_{t+1}^{inv}=f_{MoE}\left(\bm{X}_{t},\mathbf{G}_{t}\right)$
       Predict $\mathbf{G}_{t+1}^{p’}=F\mathbf{G}_{t}^{p}$ and $P_{t+1}=FP_{t}F^{T}+Q$
       Compute deviation scalar $d=Dev\left(\mathbf{T}^{ego^{\prime}},\mathbf{T}^{inv}\right)$
       Define $R^{pred}=dQ=0.2d\bar{\bm{\sigma}}$
       $\{\mathbf{G}_{m},\mathbf{R}\}=\{\mathbf{G}_{m}^{pred},R^{pred}\}$
       $hand$ := (is left-hand task ongoing) ? $lefthand$ : $righthand$
       if $Nearby(X_{t}^{root},G_{t}^{root})$ and $\neg Nearby(X_{t}^{hand},G_{t}^{hand})$ then
             $\{\mathbf{G}_{m},\mathbf{R}\}:=\{\mathbf{G}_{m},\mathbf{R}\}\cup\{\mathbf{G}_{m}^{match},R^{match}=Q\}$
       end if
      for each measurement $\{\mathbf{G}_{m}(i),\mathbf{R}(i)\}$ do
             Compute Kalman gain $\mathbf{K}=\mathbf{P}^{\prime}\left(\mathbf{P}^{\prime}+\mathbf{R}(i)\right)^{-1}$
             Update phase state $\mathbf{G}^{p}\leftarrow\mathbf{G}^{p}+\mathbf{K}\left(\mathbf{G}_{m}(i)-\mathbf{G}^{p^{\prime}}\right)$
             Update covariance matrix $\mathbf{P}\leftarrow(\mathbf{I}-\mathbf{K})\mathbf{P}$
       end for
      Update goals $\mathbf{G}_{t+1}=\{\mathbf{G}^{p},\mathbf{G}_{t+1}^{key}\}$
       $t:=t+1$
end while
ALGORITHM 1 Goal Phase Tracking with Kalman Filter
We first initialize the goal phase ${\mathbf{G}^{p}}$ and its covariance $\mathbf{P}$ using those estimated by the Kalman filter in the previous frame:
(18) $\ \ {\mathbf{G}^{p}}\leftarrow\mathbf{F}\mathbf{G}^{p},$ $\mathbf{P}\leftarrow\mathbf{F}\mathbf{P}\mathbf{F}^{T}+\mathbf{Q}$
where
(19) $\mathbf{F}$ $=\begin{bmatrix}1&0&\Delta t\\ 0&1&0\\ 0&0&1\end{bmatrix},$ $\mathbf{Q}$ $=\begin{bmatrix}\sigma_{s}^{2}&0&0\\ 0&\sigma_{f}^{2}&0\\ 0&0&\sigma_{a}^{2}\end{bmatrix},$
$\sigma_{s}^{2},\sigma_{f}^{2},\sigma_{a}^{2}$ are the process noise of the phase value, frequency, and amplitude, each of which is normalized to 0.2$\times$ of their statistical variance of the entire dataset.
Then, the goal phase ${\mathbf{G}^{p}}$ and covariance $\mathbf{P}$ are refined iteratively with the individual measurements $m=\{pred,match\}$ by its Kalman gain $\mathbf{K}_{m}$ computed using the goal phase $\mathbf{G}^{p}_{m}$ and its covariance $\mathbf{R}_{m}$:
(20) $\mathbf{G}^{p}$ $\leftarrow\mathbf{G}^{p}+\mathbf{K}_{m}\left(\mathbf{G}_{m}^{p}-\mathbf{G}^{p}\right)$
$\mathbf{P}$ $\leftarrow(\mathbf{I}-\mathbf{K}_{m})\mathbf{P}$
where $\mathbf{K}_{m}=\mathbf{P}\left(\mathbf{P}+\mathbf{R}_{m}\right)^{-1}$.
Regarding the measurement from the network, the goal phase $\mathbf{G}_{pred}^{p}$ is simply that estimated by the MoE and the covariance matrix is computed by scaling $\mathbf{Q}$ by a confidence value $c$ that is based on the bidirectional prediction constraint in Eq. 14: the more the goal-centric prediction deviates from the ego-centric prediction, the less accurate the prediction can be. Thus we compute $\mathbf{R}_{pred}$ by:
(21) $\mathbf{R}_{pred}=c\mathbf{Q}=0.2c\bar{\bm{\sigma}},$
where $c$ is computed by summing the rotational cost in Frobenius norm and the positional cost for the three key joints, as:
(22) $c$ $=Dev\left(\mathbf{T}_{t+1}^{inv},\mathbf{T}_{t+1}^{ego}\right)$
$=\sum_{j=0}^{3}\left(\alpha\|\log(\bm{r}_{j}^{ego^{T}}\bm{r}_{j}^{inv})\|_{F}+\beta\|\bm{t}_{j}^{ego}-\bm{t}_{j}^{inv}\|\right).$
where $Dev(*,*)$ computes the deviation of two 6D configurations and $\mathbf{T}_{j}^{ego}=\left[\bm{r}_{j}^{ego},\bm{t}_{j}^{ego}\right]$, $\mathbf{T}_{j}^{inv}=\left[\bm{r}_{j}^{inv},\bm{t}_{j}^{inv}\right]$ are the ego/goal-centric transformation prediction of the $j$-th key joint; the normalization parameters are set to $\alpha=1/\pi$ and $\beta$ the reciprocal of the body width. Regarding the measurement from the matched motion in the database, the corresponding phase of each frames can be precomputed by the PAE, and the covariance matrix is set to $\mathbf{Q}$.
Our Kalman filter optimally corrects the matched phase prior and simultaneously estimates a continuous-time goal trajectory in the latent space that infills the interaction keyframes, thereby enhancing the goal-driven motion quality as evaluated in § 8.3.
6. Navigation and Bimanual Planning for Pick and Placing
We now describe how we navigate the character to the target object and then schedule the bimanual manipulation for pick-and-placing. For both tasks, we compute a feature of the environment based on 3D CNN (see § 6.1) and then use it to match motions for navigation (see  § 6.2 and Fig. 2 green sub-system), where a series of locomotion data are stitched to reach the target, and bimanual scheduling (see § 6.3 and Fig. 2, Bimanual Goal Matching block and Goal Coordination block), where a series of keyframes are planned for achieving the pick-and-place actions. Fig. 6 overviews the logic of our controller in the form of a state-machine structure that produces a coordinated fully-body motion.
Figure 6. Overview of our state-machine structure for synthesis coordinate interaction-guiding goals (corresponding to Fig. 2, the bimanual goal coordination before getting the fused goal.). It adaptively allocates coordinated goals for both the hands and the entire body according to the test environment as described in § 6.3. The navigation goal before arrival (see § 6.2) and the interaction goals during manipulation (plot in green capsules) are sequentially generated to guide the DeepPhase Interaction Controller.
Figure 7. Our trajectory planner will detect the cases where the clicking target (teapot) cannot be grasped. To always achieve the interaction task, our system can sequentially remove the obstacle and then re-plan to take the target object out utilizing the available hands, and finally we can place the obstacle back to any clicking position inside the blocks. Among the trajectory visualizations, the object/skeleton from transparent to opaque sequentially shows the in-hand motion; the blue and green trajectories record the approaching and leaving motion, respectively.
6.1. Environment Features
The goal matching mechanism is to provide a motion prior for grasping/placing an object; to achieve a quick matching, we compute a volumetric scene feature around the goal position by 3D CNN. An autoencoder structure with 3 layers of convolution + max-pooling followed by 3 layers of deconvolution + upscaling is designed. A distance field of the environment pivoted at the object center obtained by fast marching is fed into the network as an input. We extracted the flattened latent code $\bm{z}\in\mathbb{R}^{16}$ together with the hand manipulation state $\bm{c}$ described in Section 4.2 as the matching feature vector from all the goal keyframes in the training set and store them in a KD tree structure.
During testing, we encode the environment around the target object $o$ to form its matching feature with the manipulation state $c$: $f=\{z_{o};\bm{c}\}$ and match $K-$nearest neighbors in the KD tree.
6.2. Navigation Matching
We now describe our scheme to guide the character in a current pose to arbitrary goal locations in the scene. Solving such a problem only by motion matching is not feasible especially if the goal location is far away. We cope with this issue by first planning a motion by 2D navigation with breadth-first search, and then finding a sequence of matched motions that can be stitched to reach the goal.
Computing the 2D Path
The 2D path from the current position of the character $\bm{p^{\prime}}$ to the goal position and direction $(\bm{p}_{g}^{root},\bm{d}_{g}^{root})$, which can be acquired by leveraging the matched root-to-furniture transformation, is computed by breadth-first search. The cost function for each position $p$ is defined as follows:
(23) $c(\bm{p})=\begin{cases}\infty,&\textit{if}\ \left(\bm{p}\right)in\ obstacles\\ angle(\bm{p}-\bm{p}_{g}^{root},\bm{d}_{g}^{root})/\pi+0.5,&\textit{if}\ ||\bm{p}-\bm{p}_{g}^{root}||_{1}<l\\ 1,&\ \textit{otherwise}\end{cases}$
where $l$ is a threshold set to $0.4m$; the second cost encourages the character to approach the goal from a direction that is parallel to the orientation of the character at the goal. The optimal path is computed using fast marching method.
Following the 2D Path by Matching
Given a 2D path computed by the breadth-first search, and the current pose of the character $\bm{p^{\prime}}$, a motion is matched from the dataset to synthesize natural locomotion that follows the path. A motion that has least deviation and has no collision with the surrounding obstacles is selected. As it will be difficult to find a match for a very long path, the path is split $3m$ from the start position and matched for each segment.
For preparing a motion dataset, navigation features defined by $\bm{N}=\{\bm{d}_{g}^{root},\bm{p}_{g}^{root},\bm{r}^{left},\bm{p}^{left},\bm{r}^{right},\bm{p}^{right},\bm{r}^{hip},\bm{p}^{hip},\bm{v}^{root}\}$ are computed for every motion between adjacent keyframes in the database, where $\bm{d}_{g}^{root}\in\mathbb{R}^{2}$ and $\bm{p}_{g}^{root}\in\mathbb{R}^{2}$ are 2D orientation/position at the goal and $\bm{r}^{keyJ}\in\mathbb{R}^{6}$, $\bm{p}^{keyJ}\in\mathbb{R}^{3}$ $\bm{v}^{root}\in\mathbb{R}^{2}$ are key joint orientation/position and root velocity relative to the root coordinate at the beginning of each motion clip. These features are used as the key to search for motion clips during navigation.
6.3. Bimanual Task Scheduler
Humans use both hands to achieve pick-and-place tasks efficiently in living environments, e.g. using one hand to open the door of the cabinet while taking out an object; to produce such movements, we propose a bimanual task scheduler that schedules the optimal series of actions by the full body.
The pipeline of the bimanual task scheduler proceeds as follows:
1. Detecting the target and auxiliary objects: The user clicks the goal object with the mouse, and then the system detects the target object (denoted as $o_{g}$) and all the other objects involved, which we call auxiliary objects $\{o_{d}\}$, such as the obstacles, and the nearby container (whose manipulable part like a door/drawer is also denoted as $o_{d}$) by ray-casting.
The system then checks if any auxiliary object is required to be manipulated. In addition to handling the closed cabinets, Fig. 7 shows another no-solution case to directly take out the inner teapot, where our system automatically removes the obstacle object before taking the target one. This is determined by whether the target object can be taken outside the cabinet without collision, based on the planned trajectory.
2. Matching the Motion for Main/Auxiliary Tasks: Next, the hands to grasp the auxiliary object and the target object are scheduled, and the poses of the three key joints, including the left hand, right hand, and the hip, are extracted from the dataset where the character grasps the auxiliary object $o_{d}$, and the target object $o_{g}$.
The scheduling is done according to the availability of the hands: If both hands are free, the goal prior matching algorithm described in Section 6.1 is conducted separately for $o_{g}$ and $o_{d}$, resulting in 4 pairs of hand-object interaction matches denoted by $\{M(H_{\{l,r\}},o_{g})$, $M(H_{\{l,r\}},o_{d})\}$ where $\{l,r\}$ represents the left and right hand. If only one hand is free, the opening and grasping tasks are allocated sequentially to the free hand; if two hands are free, the total goal costs are computed for all potential schedules, including sequentially opening and grasping with either hand and co-temporally opening and grasping with two hands. The keyframes to grasp the auxiliary object and the target object are matched such that they minimize the designed deviation cost $C$ for all possible combinations of keyframes for grasping the auxiliary object and the target object:
(24) $\min C=\min_{i_{d},i_{g}}\sum_{k=1}^{3}\{Dev(M_{k}(H_{i_{d}},o_{d}),\bm{p}^{key}_{k})$
$+wDev(M_{k}(H_{i_{d}},o_{d}),M_{k}(H_{i_{g}},o_{g}))\}$
where $M_{k}$ is the matching function that returns the 6D transformation for each key joint $k$ (left hand, right hand, and root) for the matched keyframe, $i_{d},i_{g}$ are the hand indices used for grasping the auxiliary and target objects, $\bm{p}^{key}_{k}$ is transformation of key joint $k$ in the current frame, and $w$ is a weighting set to 2.
3. Runtime Motion Synthesis: Given the series of keyframes for grasping/releasing extracted from the dataset, they are adjusted to the runtime scene by IK and used for motion synthesis. They are given to the trajectory planner to produce the trajectories of the hands and are also used as the goal poses for the DeepPhase controller. We prepare a state machine for these procedures including the 2D path planner to approach the goal as shown in Fig. 6.
7. Data Capture and Labeling
Figure 8. Snapshots of our CHOICE dataset, which incorporates different tasks like (a) cooking with ingredients and seasoning, (b) making milk tea and pouring to multiple mugs, (c) cleaning the table, and (d) other bimanual manipulations that co-temporally manipulating two objects.
Table 2. The distribution of action labels and task descriptions for the key joints. ”Cook” refers to the motion operating the pan/spatula, and ”Pour” includes both pouring drinking and the liquid ingredients.
Key Joint Action Task Seconds %
Grasp/Place 3451 12.7
Object Manipulate Cook 2845 10.5
Drink 98 0.4
Pour 3075 11.3
Door Manipulate Door opening/closing 3046 11.2
Draw Drawer Drag 1731 6.4
Idle Turn/Stand/Squat 23352 85.9
Walk Approach 3845 14.1
We extensively captured and post-processed 150 motion sequences using the Vicon Shógun  [^1] https://www.vicon.com/software/shogun/ motion capture system. The dataset comprises long-term, delicate interactions in lifelike scenes, such as a corridor with a drinking bar (Fig. 8 (d)), and a well-equipped kitchen room (Fig. 8 (a)). We approximate various involved furniture by configuring up iron frames, the plains in drawers and tables are formed by wired mesh, and all the articulated doors are glass-made. This setup ensured that every contained object was visible to the surrounding cameras. After filtering the initial data, we post-process our captures in Vicon Shógun by minimizing finger-object penetration and ensuring contact during grasping. We fit an SMPL (Loper et al., 2023) body model to the subject using SOMA (Ghorbani and Black, 2021), and export it to the Vicon Shógun body template with a hand skeleton.
Each motion sequence in our captured dataset achieves a high-level semantic goal, e.g., cooking a seasoned steak, making milk tea and drinking, or cleaning a crowded table, with an average duration of 3 minutes. The dataset exhibits substantial diversity in terms of hand movements, object types (including 14 pieces of heavy furniture, 12 freely movable objects, 6 doors, and 4 drawers) and spatial layouts, which vary across cabinets and drawers at different widths and heights, object-object relations, obstacles, and placement heights ranged from 20cm to 200cm.
In addition to the sequence-level labelling, we annotate each frame of the three key joints with action labels. The hip is labelled as either ”idle” or ”walk”, and each hand is labelled with actions like ”open a door”, ”open a drawer”, or ”approaching an object”, along with a contact label with the object. A combination of labels for the three key-joints effectively describes the current action type (see Table 2).
We further utilize the PAE’s encoder to extract the phase latent vector that implicitly represents the full-body motion information of the $[-1s,1s]$ time window pivoted on each frame. Consequently, each frame is described by its current state and the next goal state, with each state consisting of its phase latent vector, action labels, and the three key-joint transformations.
8. Experiment & Evaluation
We first describe the experimental setup for model training/testing, and then qualitatively and quantitatively evaluate our method.
8.1. Experiment Setup
Layout Data Preparation and Trajectory Planner Training
We implement our interactive character control system in Unity 3D engine and update the motion with 60 fps on a computer equipped with 5.1GHz Intel i5-13600KF and NVidia GeForce RTX 4090. For training the neural implicit trajectory planner, we reduce the network size from (Park et al., 2019) to a 6-layer MLP with 256 neurons per layer and a skip connection to the 4-th layer. We prepare 5546 interactions, each of which is composed of a pick-and-place motion and the local environment geometry represented by position/field-feature pairs § 4.2 uniformly sampled in a resolution of 2.5cm within a $(80\text{cm})^{3}$ cubic region around the target object.
Training the auto-decoder using 90% of our data takes 300 epochs by 8 hours on two Tesla A100s. At runtime, each trajectory planning will optimize the latent code by 100 steps in about 1.25 seconds.
Interaction Controller Network Training
For motion data processing, we extract 8 channels in the frequency domain with the same network structure as (Starke et al., 2022) to train the Phase Auto-encoder. Taking 120 epochs in 10 hours in all the motion data, we then pre-process all the motion data with DeepPhase features. To achieve a real-time inference over 60 fps, we designed the DeepPhase Interaction Controller network with 3 layers of 128 neurons in the Gating Network and the main network with 3 layers of 512 neurons. We train the controller network using 70% of the dataset with 150 epochs, which takes approximately 8 hours.
8.2. Trajectory Planner Evaluation
We compare our implicit neural trajectory planner with current state-of-the-art planners, including both data-driven methods and optimization-based robotic planners. We analyze the trajectory quality and evaluate the generalization capability of the planners to unseen environments.
Figure 9. Comparison of trajectory planners under representative generalization scenes with narrowed cabinets and expanded obstacles. The blue and green curves shown in two views records the approaching and leaving trajectory of the wrist, respectively. See more comparisons in our supplementary video.
8.2.1. Comparison on Motion Realism
[…]
Comparison to cuRobo: The cuRobo planner aims to produce collision-free motion with an optimal jerk cost; however, its wrist trajectory deviates from realistic human motion in two aspects. First, cuRobo tends to produce straight trajectories, where the hand and the in-hand object move too close to obstacles (see Fig. 9(b,c-2)). Also, as cuRobo is designed for rigid-body grippers, the object tends to produce minimal rotation after grasped, again resulting in little space between the object and the environment. In contrast, our planner, learned from kinematic features, employs dexterous strategies to produce human-like trajectories; for instance, when grasping a spatula out from a drawer, it makes use of the finger degrees of freedom for rotating the in-hand object while pulling it out vertically, thereby reducing the swept volume (see Fig. 9 (a-1)). Secondly, cuRobo can only produce hand trajectories with a fixed base, resulting in a lack of coordination of the locomotion with the pick-and-place task; including the base translation and orientation would be difficult with the minimum jerk criteria as human models require stepping causing many local minima. Indeed, producing human-like motion with optimization in canonical space requires a significant amount of cost/reward tuning in high dimensions, which is not practical. In contrast, our data-driven model motion can produce well-coordinated motions that are well suited for the condition represented by the latent code $\mathbf{z}$ and the condition parameters $\mathbf{c}$, eliminating the fixed hand base to achieve body-level guidance during its trajectory prediction.
Comparison to Diffusion Policy: Although learned from the same kinematic dataset, we find that the Diffusion Policy struggles to robustly produce realistic trajectories in test scenes due to the sparse dataset with high diversity. It suffers from over-straight trajectories (see Fig. 9(a,b-3) blue curves), irregular wrist orientation (see Fig. 9(b,c-3) and sudden change in height resulting in heavy placement (see  Fig. 9(d-3))
Quantitative Evaluation: The quantitative results in Table 4 also validate that our implicit neural trajectory planner can produce body motion more similar to human movement, resulting in an overall 3% lower Fréchet distance to the dataset compared with the motion following the cuRobo trajectory. Moreover, the improved motion smoothness and reduced foot sliding illustrate that our generated trajectory can be better tracked by the character, enabling it to achieve interactions with higher body coordination, benefiting from the aforementioned strengths of our kinematic data-based planner.
Figure 10. Figures from (a) to (d) sequentially show the planning of approaching a bottle with field deformation under a standard-sized shelf, 80% narrow (29.5cm in width) shelf, deep placed in the narrow shelf, and deep placed with another bottle in a narrow shelf.
Figure 11. Visualization of the implicit neural representation of the layouts and approaching trajectories uniformly sampled from the training dataset of Sec. 4.2. We project the latent vector $\mathbf{z}$ of each 3-channel field $\mathbf{D}$ to 2D by t-SNE discarding the condition labels $\mathbf{c}$. The latent space automatically clusters different types of furniture, while similar layouts are continuously distributed. The projection of all the unseen layouts (cases in grey plots) are mapped to those of similar scenes (in orange).
Table 3. Comparison among SE(3) trajectory planners for the end effector. The spare safety distance to the nearest obstacle is averaged by frame among all the collision-free successful cases, counting the pieces of trajectories within 20cm around the touching/releasing position. We run our planner (INTPs) three times at each scene with different latent initializations sampled from Gaussian noise.
Scene generalizations w/o reshaping
unsmooth. $\downarrow$ safety dist. $\uparrow$ succ. rate $\uparrow$
cuRobo 4.07 7.64 100
Diffusion 4.74 6.88 82.9
INTP-2Ch-$D_{t}$ 4.21 $\pm$ 0.10 8.15 $\pm$ 0.11 97.6
INTP-2Ch-$D_{o^{\prime}}$ 4.32 $\pm$ 0.16 7.09 $\pm$ 0.19 91.7
INTP-3Channels 4.30 $\pm$ 0.09 8.66 $\pm$ 0.13 98.8
Scene generalizations w/ reshaping & novel furniture
cuRobo 5.01 7.48 96.4
Diffusion 5.55 6.09 71.9
INTP-2Ch-$D_{t}$ 5.26 $\pm$ 0.17 7.34 $\pm$ 0.19 89.9
INTP-2Ch-$D_{o^{\prime}}$ 5.58 $\pm$ 0.29 5.74 $\pm$ 0.23 78.6
INTP-3Channels 4.85 $\pm$ 0.20 7.82 $\pm$ 0.15 95.8
8.2.2. Evaluation Under Scene Generalization
We further examine the generalization capability of our planner by applying it to unseen layouts and reshaped furniture.
Trajectory Adaptation to Scenes
Although our planner was only trained on standard-sized furniture in our motion capture scenes, it learns the spatial relationships between the trajectories and the environment, and these relationships remain consistent in novel scenes. In Fig.10, we show snapshots of the pick-and-place motions on shelves, along with the adapted trajectories in blue and the predicted time-of-arrival field $\bm{D}_{\text{toa}}$ depicted in green. Under a standard-sized shelf with an ordinary placement (Fig.10(a)), there is ample space for the right hand to reach the goal from the right side. When the shelf’s size is narrowed to 80% of the training size (see Fig.10(b)), the character instead reaches from the front and passes through the small space on the right to grasp the bottle. Further, after setting the bottle’s position deeper into the narrow shelf (Fig.10(c)), the character moves its wrist further to the left before reaching for the object. Upon adding another bottle (Fig. 10(d)), the system further adjusts the hand trajectory to avoid any collision. These experiments demonstrate that our designed three-channel fields of similar layouts can be well-mapped to a continuous latent space by the auto-decoder, enabling our planner to generalize to novel scenes.
Latent Visualization
To further illustrate the generalization capability, we analyze the latent space of our auto-decoder. Similar to how DeepSDF exhibits continuous latent representations for diverse shapes, our auto-decoder maps diverse scenes and trajectory pairs to the latent space. We project the latent codes onto a plane using t-SNE (Cieslak et al., 2020) and visualize them in Fig. 11. The projected scenes are well clustered, with unseen test scenes mapped adjacent to similar ones. This distribution reveals that even when training on scenes with various heterogeneous furniture together in one model, the auto-decoder structure effectively forms a continuous and smooth latent space for all types of scenes. This allows our runtime optimization to stably find the code that best suits the test environment and conditions, mapping it to a 6D trajectory within the distribution of plausible manipulations.
Quantitative Evaluation
To quantitatively evaluate the generalization capability of different trajectory planners, we prepare an extended test dataset composed of our test data and those from indoor scene databases  [^2] https://assetstor.unity.com/packages/3d/props/interior/kitchen-props-pack-pbr-204679; featured with various obstacle shapes, layout difference, and placement variations. Similar to Fig. 10, we synthesize unseen layouts by randomly shifting each obstacle object on/inside the involved furniture within 5cm, rotating them within ±$\pm 90^{\circ}$, and re-scaling within 0.81.25×0.8$\sim$ 1.25$\times$0.8 ∼ 1.25 ×. Also, we reshape the containers to new widths, depths, and heights ranging from 70% to 150% of their original dimensions, and remain the pick-able cases.
We compare the trajectory un-smoothness (measured by the mean acceleration), spare safety distance to obstacles from the hand and object, and the success rate among the results from our planner, the NVIDIA cuRobo planner, and the Diffusion Policy, as the examples presented in Tab. 3. Our evaluation benchmark includes 168 densely packed layouts, which will be released alongside our Unity3D project.
Our generalization benchmark reveals that the deformation of the inference field enables our planner to perform robustly in unseen, cluttered environments. In contrast, the Diffusion Policy, also as a data-driven planner, fails in narrow scenes due to collisions, as shown in cases a-3 to c-3 in Fig. 9. In these scenarios, the Diffusion model’s planning remains similar to the easier cases predominantly featured in the training data, which involve only ordinary-sized cabinets. These shortcomings become more pronounced when the layout and obstacle shapes significantly differ from the dataset (see supplementary video). Although this model could be improved by capturing more demonstrations from cabinets with greater diversity in shapes and configurations, the associated motion capture cost increases rapidly.
As shown in Table 3 part 2, our SE(3) trajectory planner achieves a high success rate over 96%, with a 4.5% higher safety distance than the cuRobo planner and 28.4% higher than the Diffusion Policy, along with significantly better motion smoothness. These results validate the benefits of our planner’s strong generalization capability and more human-like pick-and-place trajectories. The success rate of cuRobo also decreases when the scene is highly cluttered and requires more adjustments to navigate through narrow regions. In these cases, the precise geometric approximation in cuRobo cannot converge, and using a coarser geometric approximation eliminates sufficient constraints, leading to collision-prone policies.
8.2.3. Ablation Study on Distance Field Channels
We evaluate the effectiveness of our three-field setting by comparing it with a simplified two-channel auto-decoder. In these alternative models, we only keep the goal-reaching distance $D_{t}$ (as INTP-2Ch-$D_{t}$) or we merge the $D_{t}$ together with the obstacle distance $D_{o}$ into one channel $D_{o^{\prime}}$ (as INTP-2Ch-$D_{o^{\prime}}$) by the unsigned/signed distance to the nearest obstacle/target object surface, respectively. As shown in Table 3, the performance decreases significantly when using only two channels. This result confirms that having separate and complete channels for $D_{t}$ and $D_{o}$ is beneficial for high-quality trajectory predictions. Specifically, the $D_{t}$ field away from the object helps to direct the hand to approach, while $D_{o}$ field contributes to collision avoidance.
Moreover, due to the parallel optimization at runtime, reducing the number of output feature channels does not lead to a reduction in time or space complexity. Therefore, it is advantageous to utilize all accessible runtime information by distributing it across multiple channels. This approach allows our network to better infer the $D_{toa}$ field through our proposed structure.
Table 4. Ablation study and comparison to the human motion data. The unsmoothness is computed by averaging the accelerations among all joints in $cm/s^{2}$. The sliding sums up the feet’ acceleration during ground contact by $cm/s^{2}$. The rooted mean-square curvature (RMSC) of the wrist trajectory is averaged among all the approaching/leaving cases in $cm^{-1}$. The Fréchet distance evaluates the motion similarity to the dataset by extracting each 2s sliding time-windows pose as the feature vector.
[…]
w/ KF + cuRobo 6.54 7.22 691 0.0725
w/o goal phase 6.78 7.55 819 0.0982
w/ matched goal 6.36 6.80 689 0.0721
w/ predicted goal 6.30 6.86 711 0.0614
w/ KF goal 5.94 6.17 670 0.0509
Dataset 2.22 1.02 - 0.0408
Figure 12. Ablation study for our Deep Phase Interaction Controller in typical full-body interactions. We visualize the hands trace for the target approaching at (a,b), where the trajectory pieces with high curvature are highlighted by red heatmaps. In line (c) we show the bipedal trace of a pivot turning case, and line (d) shows a side-stepping case, where the skeletons from green to blue are from sequentially generated key-frames where both the feet are on the ground, with red curves highlighting the sliding pieces. We annotate each approaching trajectory with its RMS curvature in $cm^{-1}$, the mean sliding during stepping in $cm/s$, and the duration time in seconds.
8.3. Evaluation of the Pick-and-Place Motions
We show long-term interaction examples where the user produces pick-and-place motions through mouse-clicking objects in various environments. As the user clicks an object or a target location, the character automatically walks towards the object/destination and picks/places the object (see Fig. 1). The character can adapt appropriate policies to achieve the goal task with its unoccupied hand. Even under highly cluttered scenes, it executes a series of actions smoothly to remove obstacles before picking it up. Readers are referred to the supplementary video for detailed demonstrations. By testing various long-term interactions as demonstrated, we further evaluate the coordinated body motion quantitatively/qualitatively by conducting an ablation study on the goal-driven controller.
8.3.1. Ablation Study of the Deep Phase Controller
We evaluate the effectiveness of our goal-driven character controller by examining 3 variants of our DeepPhase controller, including a model that drives the character solely based on the goal key joints (w/o goal phase), a model that uses matched goal phase features obtained from the dataset (w/matched goal phase), and a model where the network is fed with goal phase features auto-regressively predicted by itself (w/ predicted goal phase). In addition to testing in the general scene described earlier (see Footnote 2), we use another test scene from our dataset composed of 20 lightweight objects and 9 pieces of furniture. A long motion sequence over 3 minutes is constructed by the user through sequential clicks on objects and furniture. We compute the smoothness, foot skating artifacts, and the Fréchet distance to the corresponding motion in the dataset (see Table 4). The results show that the use of the goal phase improves the accuracy of all measures.
Evaluation on Full-body Coordination
As the phase feature is computed based on the full body motion, it helps to coordinate all parts of the body to produce realistic movements.
When no goal phase feature is given, the legs fail to produce effective stepping to let the key joints track their goals, resulting in larger foot skating (see Tab. 4), This is evident in complex locomotion like pivot turning (see Fig.12(c-4)). Also, the upper body can struggle to recover from challenging previous poses without a goal cue, as shown in Fig.12(d-4), where the uncoordinated body causes increased foot sliding.
When the matched goal phase is used instead of that predicted by the network, the system suffers from the poor guidance to the goal. As matched goal phase are those of sparse keyframes, it cannot guide the character to conduct efficient stepping for reaching a state that is spatio-temporally apart (see Fig.12(c-3)). Also, the deviation of the current state and the matched state leads to biased directional guidance, requiring the character to take extra steps/time for reaching the goal (see Fig.12(d-3)).
Although the locomotion is natural and smooth when guided by the network-estimated goal phase, relying solely on network prediction and regression can omit some high-frequency signals, occasionally causing worse sliding performance compared to that guided by the matched goal prior (see Fig. 12(d-2)).
Since the matched goal phase provides specific goal states from similar cases, our filtered goal phase corrects the matched phase to fit the test-scene key joint goals. By combining the matched goal phase with the network prediction using the Kalman filter, we simultaneously maintain high-frequency signals and appropriately guide the locomotion to efficiently transition toward the next interaction task (see Fig. 12(c-1, d-1)).
Evaluation on Hand Control
Furthermore, the trajectory of the end-effector exhibits higher motion realism and smoothness accompanied by body coordination when using the filtered goal phase. Without the goal cue in the frequency domain, tracking only the three key joint goals spatially leads the model to focus excessively on tracking the hip or other hand goals when the manipulation takes place away from the torso. This causes large deviations when approaching the hand goals (see Fig. 12(b-4)) and results in an unsmooth trajectory of the manipulating hand.
Even with matched goal phase priors, the system can produce unsmooth wrist motions when the matched prior suddenly switches from left-hand grasping to right-hand grasping—for example, when the main hand starts to reach out for the target object after the auxiliary hand finishes opening the cabinet, as shown in Fig.12(b-3), and significant direction adjustments occur (see Fig.12(a-3)).
While the approaching motion toward the goal phase directly estimated by the MoE network slightly improves curvature results (see Fig. 12(a-2, b-2)), it still suffers from discrete switches of the goal phase. When moving the hand goal spatially along the planned approaching trajectory, the predicted goal phase updates rapidly to catch up with the goal. In case (b-2), the arm immediately stretches out to reach the target object even before the character is close enough, producing an awkward motion due to overstretching.
In contrast, the Kalman filter smoothly transitions the goal phase by combining the goal prediction with the matched grasping motion and the system’s dynamic recurrence. This results in a smooth hand trajectory well aligned with active stepping toward the target to assist the grasping process (see Fig. 12(a-1, b-1)). Additionally, it performs highly realistic small lower-body stepping during left-hand manipulation before the right-hand task, as shown in both cases (a-1, b-1). This slight body adjustment contributes to conducting the ongoing manipulation with high smoothness.
9. Discussion and Future Works
In this work, we demonstrate the effectiveness of our implicit neural trajectory planner in generating realistic pick-and-place guidance with strong generalization across various layouts. Combined with a robust goal-driven controller that simultaneously estimates the tracking state in the frequency domain, our system drives the character to perform complex interactions guided by goals of three key joints. This capability is further enhanced by our bimanual scheduler. Together, these techniques are hierarchically integrated to form the first systematic method for long-term interaction synthesis in cluttered environments. Despite these advances, our pipeline has certain limitations as follows.
Currently, our neural implicit trajectory planner is limited to pick-and-place motions. It can potentially be enhanced to handle more complex manipulation tasks, such as cooking or assembling/disassembling mechanical objects. For such purposes, it will be interesting to look into generalizing vector fields that have curls, or those that dynamically change according to the configuration of the body or the objects.
The containers we handle are limited to simple cabinets, drawers, and shelves, with fixed procedures for opening and closing them. Although this simplifies the scheduling of the planner, it cannot generalize to arbitrary containers with different types of doors, lids, and openers. Establishing a framework that can learn the structures of various container types and the procedures required to open them and retrieve contents would be an interesting research direction.
The characters are currently controlled by mouse-clicking; which is intuitive but requires some effort as the viewpoint also may need to be managed by the user. It will be interesting to look into applying LLMs to control the character by using language as an interface.
Additionally, our full-body trajectory planning for navigating towards the goal is currently based on a 2D path planner. Thus, the character cannot avoid obstacles that require the body to bend down, tilt the torso to the side or walk over. One possibility is to incorporate sensors that evaluate the distances to 3D objects for path planning and to use such sensors for motion synthesis during runtime (Holden et al., 2017; Starke et al., 2019).